OpenStudy (anonymous):

What is the Least Common Denominator of 7/3x+2/x+x/x-1? (Will post multiple choice answers if requested)

OpenStudy (campbell_st):

well the question is \[\frac{7}{3x} + \frac{2}{x} + \frac{x}{x -1}\] the 1st 2 fractions can have a common denonimator of 3x \[\frac{7}{3x} + \frac{2}{x} \times \frac{3}{3} = \frac{7 + 6}{3x}\] then the lowest common denominator of 3x and x -1 is \[\frac{13}{3x} \times \frac{x-1}{x -1} + \frac{x}{x -1} \times \frac{3x}{3x} = \frac{13(x-1)}{3x(x -1)} + \frac{x \times 3x}{3x(x-1)}\] all you need to do is simplify

OpenStudy (anonymous):

is it 3x(x-1)?

OpenStudy (campbell_st):

looks like it...

OpenStudy (anonymous):

thanks, that was one of the choices too. :D